OpenStudy (anonymous):

Evaluate (fog)(-2), given that f(x) =x^2-12x + 14 and g(x) = x^2 - 3

OpenStudy (anonymous):

To evaluate (fog)(-2), or equivalently, f(g(-2)), we must first evaluate g(-2). g(-2) = (-2)^2 - 3 = 4 - 3 = 1 Now evaluate f(g(-2)), or equivalently, f(1). f(1) = (1)^2 - 12(1) + 14 = 1 - 12 + 14 = (-11) + 14 = 3. Hence (f o g)(-2) = 3.

OpenStudy (anonymous):

Thanks